OpenStudy (anonymous):

The integral of (e^-x) cos(4x) dx.

OpenStudy (anonymous):

Use by parts twice.. keeping the original question on the left. The answer will come through recursion... does that make sense?

OpenStudy (anonymous):

Oh, so i think the reason i didn't get it is after the first "by parts" i stopped cuz it made no sense.. so you're saying if i do "by parts" again on that i should get it?

OpenStudy (anonymous):

yeah, your original "question" will show up again on the right, then you add/subract it to the other side and solve for what is left.

OpenStudy (anonymous):

Oh, true i always leave that integral out and just have the equals sign.. i should be more careful to always write what i have. similar to when i did the integral of sec^3 :P

OpenStudy (anonymous):

yes... it is very much like sec^3

OpenStudy (anonymous):

@nj1202 Are you going to show your work?

OpenStudy (anonymous):

I might, i'm just working it out on paper... so far i'm at this step: \[\int\limits_{}^{}e ^{-x}\cos(4x)dx=-e ^{-x}\cos(4x)-\int\limits_{}^{}4e ^{-x}\sin(4x)dx\]

OpenStudy (anonymous):

V = (1/4 ) sin4x

OpenStudy (anonymous):

oh... i switched my u and dv xD

OpenStudy (anonymous):

uv - ∫ vdu

OpenStudy (anonymous):

I chose the trig function to be u each time and e^-x to be dv each time... otherwise it undoes itself

OpenStudy (anonymous):

well, then that's what i did: dv = e^-x dx v = -e^-x u = cos(4x) du = -4sin(4x) dx right?

OpenStudy (anonymous):

meaning.. stay consistent... I think that you can go the other way... I just like integrating e's

OpenStudy (anonymous):

yes... that's how I did it

OpenStudy (anonymous):

minus signs... suck... watch them

OpenStudy (anonymous):

Yea i'm being careful, but i think i screwed one up cuz i have this now: \[\int\limits_{}^{}e ^{-x}\cos(4x)dx=-e ^{-x}\cos(4x)+e ^{-x}\cos(4x)+\int\limits_{}^{}e ^{-x}\cos(4x) dx\] Where's my mix up hahahahaha xD

OpenStudy (anonymous):

I know i shoud be able to add my integral to the other side, so that should be -, but does that mean the other part should be - as well?

OpenStudy (anonymous):

you are missing a 4 in the second term, which compounds to a 16 in the third term

OpenStudy (anonymous):

as well as the third term should be negative

OpenStudy (anonymous):

would you mind typing that out? it'd be of great use just to see the wrong

OpenStudy (anonymous):

oh boy... see you in a bit... :)

OpenStudy (anonymous):

hahaha xD

OpenStudy (anonymous):

\[\int\limits_{}^{}e ^{-x}\cos(4x) dx=-e ^{-x}\cos(4x)-4\int\limits_{}^{}e ^{-x}\sin(4x) dx\]for the first round... 3 negatives make the second term negative.

OpenStudy (anonymous):

that i did have :) so i was alright so far!

OpenStudy (anonymous):

u=sin(4x) du=4cos(4x) dv=e^(-x) v=-e^(-x)

OpenStudy (anonymous):

perfect oh.... forgot to multiply 4 for du... that's the 16 now i see it xD

OpenStudy (anonymous):

\[\int\limits_{}^{}e ^{-x}\cos(4x)dx=-e ^{-x}\cos(4x)+4e ^{-x}\sin(4x)-16\int\limits_{}^{}e ^{-x}\cos(4x)dx\]

OpenStudy (anonymous):

ahh now it works beautifully :)

OpenStudy (anonymous):

negative from the 4 with negative from v with negative from by parts makes a negative in the third term.

OpenStudy (anonymous):

cool... so you have divide by 17 in your answer?

OpenStudy (anonymous):

17 what? xD

OpenStudy (anonymous):

i should work this out first xD

OpenStudy (anonymous):

k

OpenStudy (anonymous):

= \[\frac{ 3e ^{-x}\cos(4x) }{ 17 }\] this it?

OpenStudy (anonymous):

\[17\int\limits_{}^{}e ^{-x}\cos(4x)dx=-e ^{-x}\cos(4x)+4e ^{-x}\sin(4x)\]

OpenStudy (anonymous):

combine like-terms and divide by 17, so i think i'm right... i'll type it in online and see.. my hw site thingy tells me if it's right or wrong.

OpenStudy (anonymous):

there is a cos term and a sin term

OpenStudy (anonymous):

oh... where'd i get the two cos xD

OpenStudy (anonymous):

got it now :D says i'm right for my hw problem!

OpenStudy (anonymous):

cool! check your signs @Chlorophyll if you factor the negative... it changes both signs inside.

OpenStudy (anonymous):

I'm getting better with this by parts stuff, just learned it thursday in lecture, so i've been doing tons of problems... this one was the first one by myself that does it twice, so it was strange. we did the sec^3 in class. but thanks euler!

OpenStudy (anonymous):

You're welcome... oddly enough I did sec^3 with a student today.

OpenStudy (anonymous):

haha xD next one is hyperbolic cosine... bring it math! :P

OpenStudy (anonymous):

@EulerGroupie Thanks, the result doesn't have negative side in front of e^-x

OpenStudy (anonymous):

I = (e^-x/ 17) ( 4sin4x - cos4x)

OpenStudy (anonymous):

yeee haaa! hyperbolics are a challenge for me because my instructors never pushed it very hard. @Chlorophyll I agree.

OpenStudy (anonymous):

haha i have never learned anything applicable with them... they just pop up to integrate xD

OpenStudy (anonymous):

I've heard that civil engineering uses them... hanging cables are apparently modeled very well by them.

OpenStudy (anonymous):

hmm that's interesting.. i almost think they can be easier than the trig functions sometimes... cuz usually it's only sinh and cosh i see, and it's always positive either der. or int. so it's less to remember xD

OpenStudy (anonymous):

Some things do seem simpler with them, and I have played with them in parametric functions along with complex numbers... what makes a hyperbola in real space makes a circle in the complex plane... its very wierd. Anyway, you may be on your own with the next one... it is after midnight here.

OpenStudy (anonymous):

haha yea it's 2 am where i'm at haha xD i work more efficiently in late hours... and i can sleep in xD no classes on sunday :D

OpenStudy (anonymous):

Good luck! See ya' around.

OpenStudy (anonymous):

Thanks buddy :D